In a random sample of 1,000 people, it is found that 8.7% have a liver ailment. Of those who have a liver ailment, 6% are heavy drinkers, 60% are moderate drinkers, and 34% are nondrinkers. Of those who do not have a liver ailment, 8% are heavy drinkers, 43% are moderate drinkers, and 49% are nondrinkers.
If a person is chosen at random, and he or she is a heavy drinker, what is the empirical probability of that person having a liver ailment?

Answer by stanbon(75887)

In a random sample of 1,000 people, it is found that 8.7% have a liver ailment.
--------
Of those who have a liver ailment, 6% are heavy drinkers, 60% are moderate drinkers, and 34% are nondrinkers.
--------------------------------
Of those who do not have a liver ailment, 8% are heavy drinkers, 43% are moderate drinkers, and 49% are nondrinkers.
If a person is chosen at random, and he or she is a heavy drinker, what is the empirical probability of that person having a liver ailment?
----
P(liver ailment | heavy drinker) = 0.087*0.08 = 0.00696
---
Ans: # = 0.00696*1000 = 6.96 or approximately 7
------------
Cheers,
Stan H.
-----------